OpenStudy (anonymous):

if (/) is an acute of a right triangle and cos (/) is 5/7 find the values of the other five trig functions of (/)

OpenStudy (luigi0210):

Let us draw! :P

OpenStudy (luigi0210):

do you know how to draw that?

OpenStudy (anonymous):

that i am trying to do know, with minimal luck

OpenStudy (luigi0210):

do you know what cos is?

OpenStudy (anonymous):

yes 5/7 & secant is 7/5

OpenStudy (luigi0210):

exactly! it is also:\[\cos=\frac{ adj }{ hyp }=\frac{ 5 }{ 7 }\]

OpenStudy (luigi0210):



OpenStudy (luigi0210):

can you find the missing side now?

OpenStudy (anonymous):

honest on brain freeze maybe because i got sin x/5, & lost stuck to next step

OpenStudy (anonymous):

i'm trying a^2+b^2=c^2, but i think that step is wrong

OpenStudy (luigi0210):

No you are right!

OpenStudy (luigi0210):